# Spontaneous orbital selective Mott phase in the two band Hubbard model

### v1

@@ IV Discussions
Studying the two-band Hubbard model with strong on-site repulsion, we explored a spontaneous orbital-selective phase in the vicinity of half-filling for weak to moderate interorbital hopping. Despite the stability of a sizeable spontaneous orbital symmetry breaking phase for finite interorbital hopping, the hybridization between the itinerant and localized orbitals remains nonvanishing. This finding aligns with results obtained using single-site DMFT [26] and can be understood as the inevitability of electron transfer between the two orbitals, leading to finite hybridization.
@@
Therefore, our analysis demonstrates that orbital-selective physics does not necessarily exhibit Kondo breakdown transitions. This result does not contradict previous works that interpret Kondo breakdown as a selective Mott transition [40]. In that study [40], the model involves free $c$-electrons interacting with strongly correlated $f$-electrons through inter-orbital hopping and inter-orbital Coulomb repulsion, leading to a complete breakdown phase. Although we do not find a complete Kondo breakdown phase within our model, hybridization is only finite where the two bands cross below the Fermi energy. Furthermore, significant orbital differentiation and a small Fermi surface remain across a range of interorbital hopping values. We confirmed the stability of our results against a non-zero inter-orbital repulsion $U^{\prime}$ when it is a perturbation on par with the hopping. While some proposals suggest the possibility of a complete Kondo breakdown for large $U^{\prime}$ [41] is currently beyond the scope of this study. and remains an open question for the future.
@@
The Fermi volume jumps sharply at the transition between the OSMP and OUP phase. In OSMP, one of the bands spontaneously flattens and goes completely below the Fermi energy, thereby retaining the overall orbital character [51]. The presence of a filled band requires the other orbital to go below average density to maintain the total density. Consequently, a jump in the Fermi volume occurs when both bands are hybridized as the system transitions to orbitally uniform density. Our finding indicates that any detection of a small Fermi surface in experiments is not equivalent to a complete breakdown of hybridization between orbitals.
@@
The idea of a critical Fermi surface [52] where the quasiparticle weight of one of the orbitals vanishes continuously does not apply to the abrupt OSMP transition observed in this study. A critical Fermi surface lacks well defined excitations due to the vanishing quasiparticle residue. Here, the transition is discontinuous, as evidenced by the sharp jumps in several quantities, and therefore, it lacks a quantum critical point. The entire Fermi surface abruptly disappears at the OSMP transition, and the quasiparticle residue of one orbital vanishes at the transition.
@@
This work highlights the possibility of stabilizing a spontaneous orbital selective phase within a two-band model and only on-site Hubbard repulsion. Previous studies on OSMP focus on explicit symmetry breaking by tuning the crystal field splitting of the orbitals [53, 54, 55] or varying the intra-band Coulomb repulsion [56]. Other studies found spontaneous OSMP [57] with a negative Hund’s coupling $J_{H}$ in a three orbitals model. The OSMP transition is reminiscent of a discontinuous transition, with a large parameter region where the OSMP phase is metastable.
@@
This study focuses on the paramagnetic solution of the two-band Hubbard model. The localized orbitals in the OSMP, acting as quenched spins, are expected to order antiferromagnetically, similar to a single-band Hubbard model below the Néel temperature  [58, 59]. An intriguing future direction would be investigating this orbital-selective phase in a lattice with frustrated spins, such as a triangular lattice, where magnetic order is suppressed. For an odd number of localized electrons per unit cell, the resulting paramagnetic spin liquid exhibits topological order [60], featuring fractionalized spinon excitations coexisting with conduction electrons from the nonlocalized orbital [51].
@@ V Acknowledgement
CP and AB thanks funding from CEFIPRA. All computations are performed on the IPhT Kanta cluster.
@@ Appendix A Roth decoupling
This scheme begins with the provided composite basis $\mathbf{\Psi}$. We call $\mathcal{B}_{i\alpha}=\{B^{1}_{i\alpha},B_{i\alpha}^{2},\cdots,B_{i\alpha}^{N}\}$ the set of bilinear fermion operators at a site $i$ for the $\alpha$ orbital and we suppress the spin index. For instance $B^{1}_{i\alpha}=c^{\dagger}_{i\alpha}c_{i\alpha}$ is the local density operator. Owing to the fermionic commutation relations, the Roth decoupling scheme can be applied independently to $\langle B^{n_{1}}_{ix}B^{n_{2}}_{jx}\rangle$, $\langle B^{n_{1}}_{ix}B^{n_{2}}_{jy}\rangle$ and $\langle B^{n_{1}}_{iy}B^{n_{2}}_{jy}\rangle$ which means, in practice, there will be three independent sets of self-consistent equations for the $xx$, $xy$ and $yy$ sector. The solutions of these equations provide access to the 2-point correlation functions of each sector.
Roth decoupling allows us to find a set of self-consistent equations between the following static correlators $\langle B_{i\alpha}^{n}B_{j\sigma}^{m}\rangle$, $\forall n,m\in{1,\cdots,N}$ with $i$. Defining $\mathbf{c}_{i}=(c_{i\tau\uparrow},c_{i\tau\downarrow},c_{i\tau\uparrow}^{\dagger},c_{i\tau\downarrow}^{\dagger})$ the fermionic basis at each site $i$ for each orbital $\tau$. then $\forall n$, we can express $B^{n}_{i\tau}$ as $B^{n}_{i\tau}=\mathbf{c}_{ia(n)}\mathbf{c}_{ib(n)}$ with two indices $a(n)$ and $b(n)$ that depend on $n$.
@@
We can introduce the two-particle static correlation function vector in each orbital sector
@@ B.1 Energy per-site
The first term is the intraorbital kinetic energy $\langle\mathcal{H}_{t}\rangle$ and the second term is interorbital kinetic energy $\langle\mathcal{H}_{\lambda}\rangle$. The third term is the potential energy $\langle\mathcal{H}_{U}\rangle$. We evaluate in terms of the correlation function
@@ B.2 Spectral function and density of states
with $\psi(\alpha),\psi^{\prime}(\beta)$ representing a summation over the composite operator of orbital character $\alpha\beta$. The spectral function is given by
@@
The total intraorbital spectral function is given by ${A(\mathbf{k},\omega)=\sum_{\alpha}A_{\alpha\alpha}(\mathbf{k},\omega)}$. Similarly the interorbital spectral function is given by ${A_{\delta}(\mathbf{k},\omega)=\sum_{\alpha}A_{\alpha\bar{\alpha}}(\mathbf{k},\omega)}$. The orbital resolved density of states can be evaluated as,
@@
Figure 8: Momentum resolved electron density along the $\Gamma$ to M high symmetry line. The electron density is fixed at $n=1.7$ and the blue curve corresponds to the OSMP while the two other are in the ’Fermi liquid’ phase. In the vicinity of Fermi momentum a jump is observable. Since the OUP phase exhibits two electron pockets, two distinct jumps are observable. The OSMP only presents one jump, but the density does not vanish at M due to the fully filled orbital. According to Fermi liquid theory the values of these jumps are the quasiparticle weights. On 5 (A) we report the value of the jump as dots, and confirm it matches the other definition of the quasiparticle weight.
@@ B.3 QuasiParticle-Weight $Z$
with $\psi(\alpha),\psi^{\prime}(\beta)$ representing a summation over the composite operator of orbital character $\alpha\beta$. This can be rewritten as
@@
In Fig. 8, $N_{k}$ is plotted along the $\Gamma$ to M line for several $\lambda$ at $n=1.70$. In the orbitally uniform phase phase, two jumps are observed due to the presence of two electron pockets in the Fermi surface. The quasiparticle weight is equal to the sum of the two jumps. The OSMP only exhibits one jump due to the disappearance of one of the pockets, but has a residual electron density at M due to the fully filled band. On Fig. 5 we report the value of the jump as colored dots. The dots are indeed fitting the continuous line representing the quasiparticle weight, proving the correspondence between the two methods.
Appendix C OSMP phase at $\lambda=0$
Figure 9: Energy per site for the single-band Hubbard model as a function of electron density $n$. The red diamond, green square, and orange circle indicate the density and energy at $n=1.0$, $n=0.9$, and $n=0.8$ respectively. For two orbitals, if the the total electron density is $n=1.8$. Without interorbital hopping this demonstrates that it is energetically favorable to have one orbital at $n=1$ and the other at $n=0.8$ (OSMP phase), rather than both orbitals at $n=0.9$ (OUP phase).
The hole doping $\delta$ for each orbital in two-orbital systems presents two possibilities. One option is for the system to have a uniform density of holes, such that $n=1-\delta$, as in the OUP phase. Alternatively, it may be energetically favorable for the holes to concentrate in a single orbital, resulting in a density of $n=1-2\delta$, while the other orbital remains half-filled at $n=1$.
When $\lambda=0$, the two orbitals decouple entirely from each other. The stability of the OSMP or OUP phase can then be understood by studying the energy per site $E_{s}$ of the single-band Hubbard model. Fig. (9) shows the energy evolution per site with electron density. There is a notable decrease in energy close to half-filling, indicating the proximity to the Mott phase. Such energy reduction near half-filling makes it energetically favorable to have one layer with a density of $n=1$ and the other layer with a density of $n=1-2\delta$ rather than having both layers with a density of $n=1-\delta$. Note that OSMP is favored up to critical total density $n_{c}$ around half-filling, beyond which the system becomes orbitally uniform.

### v2

@@ IV Discussions
Studying the bilayer Hubbard model with strong on-site repulsion, we explored a spontaneous layer-selective phase in the vicinity of half-filling for weak to moderate inter-layer hopping. Despite the stability of a sizeable spontaneous layer symmetry broken phase for finite inter-layer hopping, the hybridization between the itinerant and localized layer remain nonvanishing. This finding aligns with results obtained using single-site DMFT for selective Mott phase [26] and can be understood as the inevitability of electron transfer between the two layers(or orbitals), leading to finite hybridization.
@@
Therefore, our analysis demonstrates that selective Mott physics does not necessarily exhibit full Kondo breakdown transitions. This result does not contradict previous works that interpret Kondo breakdown as a selective Mott transition [47]. In that study [47], the model involves free $c$-electrons interacting with strongly correlated $f$-electrons through inter-orbital hopping and inter-orbital Coulomb repulsion, leading to a complete breakdown phase. Although we do not find a complete Kondo breakdown phase within our model, hybridization is only finite where the two bands cross below the Fermi energy. Furthermore, significant layer differentiation and a small Fermi surface remain across a range of inter-layer hopping values. We confirmed the stability of our results against a non-zero inter-layer repulsion $U^{\prime}$ when it is a perturbation. While some proposals suggest the possibility of a complete Kondo breakdown for large $U^{\prime}$ [48] is currently beyond the scope of this study. and remains an open question for the future.
@@
The Fermi volume jumps sharply at the transition between the LSMP and LUP phase. In LSMP, one of the bands spontaneously flattens and goes completely below the Fermi energy, thereby retaining the overall layer character [63]. The presence of a filled band requires the other band to go below average density to maintain the total density. Consequently, a jump in the Fermi volume occurs when both bands are hybridized as the system transitions to layer uniform phase. Our finding indicates that any detection of a small Fermi surface in experiments is not equivalent to a complete breakdown of hybridization between two degrees of freedom.
@@
The idea of a critical Fermi surface [64] where the quasiparticle weight of one of the layers vanishes continuously does not apply to the abrupt LSMP transition observed in this study. A critical Fermi surface lacks well defined excitations due to the vanishing quasiparticle residue. Here, the transition is discontinuous, as evidenced by the sharp jumps in several quantities, and therefore, it lacks a quantum critical point. The entire Fermi surface abruptly disappears at the LSMP transition, and the quasiparticle residue of one layer vanishes at the transition.
@@
This work highlights the possibility of stabilizing a spontaneous Mott selective phase within a bilayer system and only on-site Hubbard repulsion. Studies on selective Mott transitions focus on explicit symmetry breaking by tuning the crystal field splitting of the layers [65, 66, 67] or varying the intra-band Coulomb repulsion [68]. Other studies found spontaneous OSMP [69] with a negative Hund’s coupling $J_{H}$ in a three orbitals model. The LSMP transition is reminiscent of a discontinuous transition, with a large parameter region where the LSMP phase is metastable.
@@
This study focuses on the paramagnetic solution of the bilayer Hubbard model. The localized layer in the LSMP, acting as quenched spins, are expected to order antiferromagnetically, similar to a single-band Hubbard model below the Néel temperature  [70, 71]. An intriguing future direction would be investigating this selective Mott phase in a lattice with frustrated spins, such as a triangular lattice, where magnetic order is suppressed. For an odd number of localized electrons per unit cell, the resulting paramagnetic spin liquid exhibits topological order [72], featuring fractionalized spinon excitations coexisting with conduction electrons from the nonlocalized layer [63]. Recently, many bilayer twisted materials have been shown to exhibit correlated behaviors analogous to Mott physics [73, 74, 75]. Extending the current study to twisted bilayer models is therefore a promising avenue to explore layer selective Mott phase in flat bands.
@@ V Acknowledgement
CP and AB thanks funding from CEFIPRA (Grant No. 6704-3). All computations are performed on the IPhT Kanta cluster.
@@ Appendix A Roth decoupling
This scheme begins with the provided composite basis $\mathbf{\Psi}$. We call $\mathcal{B}_{i\alpha}=\{B^{1}_{i\alpha},B_{i\alpha}^{2},\cdots,B_{i\alpha}^{N}\}$ the set of bilinear fermion operators at a site $i$ for the $\alpha$ layer and we suppress the spin index. For instance $B^{1}_{i\alpha}=c^{\dagger}_{i\alpha}c_{i\alpha}$ is the local density operator. Owing to the fermionic commutation relations, the Roth decoupling scheme can be applied independently to $\langle B^{n_{1}}_{ix}B^{n_{2}}_{jx}\rangle$, $\langle B^{n_{1}}_{ix}B^{n_{2}}_{jy}\rangle$ and $\langle B^{n_{1}}_{iy}B^{n_{2}}_{jy}\rangle$ which means, in practice, there will be three independent sets of self-consistent equations for the $xx$, $xy$ and $yy$ sector. The solutions of these equations provide access to the 2-point correlation functions of each sector.
Roth decoupling allows us to find a set of self-consistent equations between the following static correlators $\langle B_{i\alpha}^{n}B_{j\sigma}^{m}\rangle$, $\forall n,m\in{1,\cdots,N}$ with $i$. Defining $\mathbf{c}_{i}=(c_{i\tau\uparrow},c_{i\tau\downarrow},c_{i\tau\uparrow}^{\dagger},c_{i\tau\downarrow}^{\dagger})$ the fermionic basis at each site $i$ for each layer $\tau$. then $\forall n$, we can express $B^{n}_{i\tau}$ as $B^{n}_{i\tau}=\mathbf{c}_{ia(n)}\mathbf{c}_{ib(n)}$ with two indices $a(n)$ and $b(n)$ that depend on $n$.
@@
We can introduce the two-particle static correlation function vector in each layer sector
@@ B.1 Energy per-site
The first term is the intralayer kinetic energy $\langle\mathcal{H}_{t}\rangle$ and the second term is inter-layer kinetic energy $\langle\mathcal{H}_{\lambda}\rangle$. The third term is the potential energy $\langle\mathcal{H}_{U}\rangle$. We evaluate in terms of the correlation function
@@ B.2 Spectral function and density of states
with $\psi(\alpha),\psi^{\prime}(\beta)$ representing a summation over the composite operator of layer character $\alpha\beta$. The spectral function is given by
@@
The total intralayer spectral function is given by ${A(\mathbf{k},\omega)=\sum_{\alpha}A_{\alpha\alpha}(\mathbf{k},\omega)}$. Similarly the inter-layer spectral function is given by ${A_{\delta}(\mathbf{k},\omega)=\sum_{\alpha}A_{\alpha\bar{\alpha}}(\mathbf{k},\omega)}$. The layer resolved density of states can be evaluated as,
@@
Figure 9: Momentum resolved electron density along the $\Gamma$ to M high symmetry line. The electron density is fixed at $n=1.7$ and the blue curve corresponds to the LSMP while the two other are in the ’Fermi liquid’ phase. In the vicinity of Fermi momentum a jump is observable. Since the LUP phase exhibits two electron pockets, two distinct jumps are observable. The LSMP only presents one jump, but the density does not vanish at M due to the fully filled layer. According to Fermi liquid theory the values of these jumps are the quasiparticle weights. On 6 (A) we report the value of the jump as dots, and confirm it matches the other definition of the quasiparticle weight.
@@ B.3 QuasiParticle-Weight $Z$
with $\psi(\alpha),\psi^{\prime}(\beta)$ representing a summation over the composite operator of layer character $\alpha\beta$. This can be rewritten as
@@
In Fig. 9, $N_{k}$ is plotted along the $\Gamma$ to M line for several $\lambda$ at $n=1.70$.
Appendix C LSMP for different inter-layer hopping symmetry
For systems with different interlayer orbital symmetry direct interlayer-site hopping is usually forbidden due to symmetry constraints between different orbitals. As a result, inter-layer nearest-neighbor hopping becomes the dominant term, which can destabilize the layer-selective Mott phase [23]. In this appendix, we examine both extended $s$- and $d$-wave inter-layer hopping and demonstrate that it produces similar physical behavior as on-site inter-layer hopping. The extended $s$- and $d$-wave interlayer hopping read respectively in moemtum space as
$\mathcal{H}_{\lambda}^{s}(\mathbf{k})=-2\lambda\sum\limits_{\mathbf{k}\sigma}\left(\cos(k_{x})+\cos(k_{y})\right)\left(c^{\dagger}_{\mathbf{k}x\sigma}c_{\mathbf{k}y\sigma}+h.c.\right)$ (82)
$\mathcal{H}_{\lambda}^{d}(\mathbf{k})=-2\lambda\sum\limits_{\mathbf{k}\sigma}\left(\cos(k_{x})-\cos(k_{y})\right)\left(c^{\dagger}_{\mathbf{k}x\sigma}c_{\mathbf{k}y\sigma}+h.c.\right)$ (83)
Figure 10: (A) Disparity in layer densities ($\delta n$) as a function of $\lambda$ at $n=1.8$ and $U=20t$ for both extended $s$-wave (blue curve) and $d$-wave (orange curve) nearest-neighbor hopping. (B) Compares the energy per site between the LUP and LSMP phase for both $s$-wave (blue curve) and $d$-wave (orange curve) nearest-neighbor hopping. Dashed lines are for the LUP solutions and thick lines for the LSMP ones.
In Fig. (10a), we show that for a fixed $U=20t$, we observe a transition from the LSMP to LUP phase at low doping. Similar to the direct on-site interlayer hopping the LSMP is energetically favorable for small $\lambda$ as shown in Fig. (10 b) for both different form factor of the nearest neighbor hopping.